40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5367     Entry time: Thu Sep 8 20:13:24 2011
Author: kiwamu 
Type: Update 
Category: IOO 
Subject: MC is back to full power 

[Suresh / Kiwamu]

 The attenuator was removed and now the MC is happily locked with the full power of 1.2 W.

 

(what we did)

 + replaced the perfect reflector, which was before the MCREFL_PD, by a 10% beam splitter like it used to be.

 + removed the attenuator (combination of HWP and PBS).

 + realigned the beam path on the AP table, including the MCREFL path and WFS path.

 + made the aperture of the MC2F camera narrower in order to avoid a saturation.

 + aligned the MC suspensions so that it resonates with the TEM00 mode.

 + put a ND filter on the AS camera

 

(notes)

C1:IOO-MC_RFPD_DCMON = 0.98 (locked)

C1:IOO-MC_TRANS_SUM = 17500 (locekd)

 

(next things to do)

 + measurement of the spot positions on each MC mirror.

 + centering of the beam spot by steering the input mirrors on the PSL table

ELOG V3.1.3-